OpenStudy (anonymous):

What is the volume of a square prism whose base edges are 4 inches and height is 12 inches? 252 in3 576 in3 48 in3 192 in3

OpenStudy (anonymous):

isnt it 192 in3?

OpenStudy (anonymous):

4x4x12=192

OpenStudy (campbell_st):

Volume of a Prism is V = Area * height The area of the square is A=4^2 = 16 V = 16 * 12